0.085 m3 of air are compressed isentropically from 1937 kPa and -4 oC to 5873 kPa . Find (a) t2 in oC and V2, (b) Wnf, and © H and U.


1
Expert's answer
2021-12-03T10:59:59-0500

"T_2=T_1( \frac{P_2}{P_1} )^ \frac{k-1}{k}"

Where:

"T_2=" final temperature

"T_1=269K"

"P_2=5873kPa"

"P_1=1937kPa"

"k=1.4=" constant

"T_2=269×( \frac{5873}{1937} )^ \frac{1.4-1}{1.4} =369.3K=96.3^0C"

Also:

"\frac{T_2}{T_1} =(\frac{V_1}{V_2} )^ {k-1}"

"\frac{369.3}{269}=(\frac{0.085}{V_2})^{(1.4-1)}"

"V_2=0.0385M^3"